OpenStudy (anonymous):

A line passes through the point (-3,5) and has a slope of -5 . Write an equation for this line.

OpenStudy (anonymous):

Use Point Slope Form: \[\large y - y_1 = m(x - x_1)\] Here : \(m = -5, y_1 = 5, x_1 = -3\) Just plug in the values and solve..

OpenStudy (anonymous):

y-5=-5(x+3)

OpenStudy (anonymous):

Yep.. Now solve Right hand side more..

OpenStudy (anonymous):

Distribute -5 to the brackets..

OpenStudy (anonymous):

y-5=-5x+15

OpenStudy (anonymous):

y-5=-5x-15

OpenStudy (anonymous):

Yep, Good going.. Now add 5 both the sides..

OpenStudy (anonymous):

y=-5x-10?

OpenStudy (anonymous):

Yep...